40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Jul 30 09:13:14 2020, shruti, DailyProgress, Lab Work, beam profile prep 
    Reply  Fri Jul 31 08:51:00 2020, shruti, DailyProgress, Lab Work, beam profile prep 
Message ID: 2546     Entry time: Thu Jul 30 09:13:14 2020     Reply to this: 2547
Author: shruti 
Type: DailyProgress 
Category: Lab Work 
Subject: beam profile prep 

Notice of lab entry: I will be at the cryo lab today (30 Jul 20) 9am-5pm.

Purpose: Beam profile the Rio planex laser

- Turned on the laser that goes into the east cavity on. Turned on the Tenma supply, then the laser current driver  D1500207 labeled'E'. Saw a tiny, but bright, green spot on the detector card.

- Located razor blade and translation stage, for the beam profile measurement but not sure which power meter/ photodiode to use for 1550 nm. I plan to move the optics on the path of the east laser (to the east cavity) in order to do this measurement and later set up the initial ring cavity

- Aaron suggested I do another airflow measurement within the lab. Without the scrubber, there seems to be no measurable flow around either the Cryo-Q table or the cantilever table. I moved the scrubber into the lab near the workspace at the north side of the lab and let it run on medium speed. A few cm from the scrubber it the airflow rate is almost 5000 CFM, but it becomes <40 CFM even as close as the Cryo-Q table and again becomes negligible near the cantilever table.

- I set up the translation stage, moved relevant optics, but will continue to measurement later after I've researched and located a 1550 nm power meter

Two of the four overhead lights over the west (cantilever) table don't work anymore. Three other overhead lights are very dim/out.

 

 

ELOG V3.1.3-